the age of man is 10 years more than three times his brother's age. Find the age if 5 years from now he will be 4 years more than twice hi brothers age.
Answer by ptaylor(2198)

Let x=brother's age now; x+5=age 5 years from now
Then 3x+10=age of the man now; 3x+10+5=3x+15=age 5 years from now

And we are told that:
3x+15=2(x+5)+4
3x+15=2x+10+4
3x+15=2x+14 subtract 15 and also 2x from each side
3x-2x=-15+14
x=-1 WHOAAAAAAAA!!! CAN'T HAVE NEGATIVE AGES
THE FACT OF THE MATTER IS THAT IN 5 YEARS, THE MAN WILL BE 3(X+5) OR THREE TIMES AS OLD AS HIS BROTHER (X+5)
Hope this helps---ptaylor